If the price of a product is given by p=100e^q/(q+1), where q represents the number of hundreds of units, what will be the price when the producers are willing to supply 300 units? Hint: e~2.718
Answer by RAY100(1637)

P=100e^(q/(q+1))
where q=300/100=3
P=100e^(3/4)=100e^.75=100*2.117=211.70